1) find the area of a trapezoid which has a measure of 7 meters for its height and 4 meters and 6 meters for its two bases
Answer by checkley71(8403)

AREA=(B1+B2)/2*H
AREA=(4+6)/2*7
AREA=10/2*7
AREA=5*7=35M^2